# From Interpolating Formulas to Separating Languages and Back Again

### v1

@@ 6 Back Again: Interpolant Existence in Linear Temporal Logic
On the other hand, Theorems 5.38 and 6.46 suggest a characterisation of regular languages in terms of existentially quantified $\mathsf{LTL}$-formulas. More precisely, define $\exists\mathsf{LTL}$ as the language that comprises $\mathsf{LTL}$-formulas possibly prefixed by quantifiers $\exists q$ over propositional variables $q$ that are interpreted by taking $\mathfrak{M},n\models\exists q\,\varphi$ iff $\mathfrak{M}^{\prime},n\models\varphi$, for some $\mathfrak{M}^{\prime}$ that may only differ from $\mathfrak{M}$ on the interpretation of $q$. Observe that the truth-condition for $\exists q_{1}\dots\exists q_{n}\varphi$ can be expressed by an $\mathsf{MSO}$-formula of the form $\exists Q_{1}\dots\exists Q_{n}\,\psi_{\varphi}$, where $\psi_{\varphi}$ is the standard translation (in the signature $\sigma_{\textit{lin}}$) for $\varphi$ and the $Q_{i}$ are fresh unary predicate symbols (not elements of $\sigma_{\textit{lin}}$). Moreover, the following is a consequence of Theorems 5.38 and 6.46:
@@
and so $\varphi\models_{\mathsf{LTL}}\psi$. Any interpolant $\chi$ for $\varphi,\psi$ in $\mathsf{LTL}$ should be variable-free and express that the number of points in an $\emptyset$-model is even; the standard translation of such a $\chi$ should be an $\mathsf{FO}(<)$-sentence expressing parity of the length of a finite strict linear order, which is impossible by (2).
So far, the decidability status of the IEP for $\mathsf{LTL}$ has been open, with the methods of analysing the problem discussed in Section 3 being non-applicable in this case. The next theorem settles this open question by employing Theorem 6.46 to reduce—for the price of a multi-exponential blow-up—the IEP for $\mathsf{LTL}$ to $\mathsf{FO}(<)$-separability of regular languages and then using Theorems 5.39, 5.41 and Corollary 5.42.
@@
The IEP for $\mathsf{LTL}$ is decidable in $4\textsc{ExpTime}$, and PSpace-hard.
@@
As for the computational complexity of deciding the IEP for $\mathsf{LTL}$, the above reduction gives a $4\textsc{ExpTime}$ upper bound: the size of an NFA $\mathfrak{A}$ specifying the language defined by an $\mathsf{LTL}$-formula $\varphi$ is exponential in the size of $\varphi$ [113], the NFA specifying ${\boldsymbol{L}}_{\varphi}$ is a projection of $\mathfrak{A}$ [99], and so has size linear in the size of $\mathfrak{A}$; another exponential is to turn this NFA to a DFA, and then two more exponentials come from Corollary 5.42.
@@
$\mathsf{LTL}$ over $(\omega,<)$. $\mathsf{LTL}$-formulas can also be interpreted in infinite temporal models of the form $(\omega,<,\mathfrak{v})$. We denote the resulting logic by $\mathsf{LTL}_{\omega}$. This kind of temporal models can also be regarded as $\sigma_{\textit{lin}}$-structures of the $\mathsf{FO}$-signature $\sigma_{\textit{lin}}$ above. As discussed above, the analogues of Theorems 5.39 and 5.41 hold for $\omega$-languages. Moreover, Theorem 6.46 also holds in this setting, and so a proof similar to that of Theorem 6.50 shows that the IEP for $\mathsf{LTL}_{\omega}$ is also decidable, with its exact complexity remaining also open.
@@
It is also of interest to note that quantified $\mathsf{LTL}$ is the smallest ‘natural’ logic containing $\mathsf{LTL}[\Diamond]$ and having the CIP [47].

### v2

@@ 6 Back Again: Interpolant Existence in Linear Temporal Logic
On the other hand, Theorems 5.38 and 6.46 suggest a characterisation of regular languages in terms of existentially quantified $\mathsf{LTL}$-formulas. More precisely, define $\exists\mathsf{LTL}$ as the language that comprises $\mathsf{LTL}$-formulas possibly prefixed by quantifiers $\exists q$ over propositional variables $q$ that are interpreted by taking $\mathfrak{M},n\models\exists q\,\varphi$ iff $\mathfrak{M}^{\prime},n\models\varphi$, for some $\mathfrak{M}^{\prime}$ that may only differ from $\mathfrak{M}$ on the interpretation of $q$. Observe that the truth-condition for $\exists q_{1}\dots\exists q_{n}\varphi$ can be expressed by an $\mathsf{MSO}$-formula of the form $\exists Q_{1}\dots\exists Q_{n}\,\psi_{\varphi}$, where $\psi_{\varphi}$ is the standard translation (in the signature $\sigma_{\textit{lin}}$) for $\varphi$, and each $Q_{i}$ is the unary predicate symbol corresponding to $q_{i}$. Moreover, the following is a consequence of Theorems 5.38 and 6.46:
@@
[…]
So far, the decidability status of the IEP for $\mathsf{LTL}$ has been open, with the methods of analysing the problem discussed in Section 3 being non-applicable in this case. The next theorem settles this open question by employing Theorem 6.46 to reduce—for the price of a multi-exponential blow-up—the IEP for $\mathsf{LTL}$ to $\mathsf{FO}(<)$-separability of regular languages and then using Theorems 5.39, 5.41 and 5.42.
@@
The IEP for $\mathsf{LTL}$ is decidable in $3\textsc{ExpTime}$, being PSpace-hard.
@@
As for the computational complexity of deciding the IEP for $\mathsf{LTL}$, the above reduction gives a $3\textsc{ExpTime}$ upper bound: the size of an NFA $\mathfrak{A}$ specifying the language defined by an $\mathsf{LTL}$-formula $\varphi$ is exponential in the size of $\varphi$ [113], the NFA specifying ${\boldsymbol{L}}_{\varphi}$ is a projection of $\mathfrak{A}$ [99], and so has size linear in the size of $\mathfrak{A}$; and then two more exponentials come from Theorem 5.42.
@@
$\mathsf{LTL}$ over $(\omega,<)$. $\mathsf{LTL}$-formulas can also be interpreted in infinite temporal models of the form $(\omega,<,\mathfrak{v})$. We denote the resulting logic by $\mathsf{LTL}_{\omega}$. This kind of temporal models can also be regarded as $\sigma_{\textit{lin}}$-structures of the $\mathsf{FO}$-signature $\sigma_{\textit{lin}}$ above. As discussed above, the analogues of Theorems 5.39, 5.41, and 5.42 hold for $\omega$-languages. Moreover, Theorem 6.46 also holds in this setting, and so a proof similar to that of Theorem 6.50 shows that the IEP for $\mathsf{LTL}_{\omega}$ is also decidable, with its exact complexity remaining also open.
@@
It is also of interest to note that quantified $\mathsf{LTL}$ (with both existential and universal quantification over propositional variables) is the smallest ‘natural’ logic containing $\mathsf{LTL}[\Diamond]$ and having the CIP [47].